OpenStudy (anonymous):

Which exponential function goes through the points (1, 8) and (4, 512)?

OpenStudy (solomonzelman):

okay, so an exponential function is: \(\Large\color{black}{ y=ab^x }\) (where \(\large\color{black}{ a }\) and \(\large\color{black}{b }\) are constants)

OpenStudy (solomonzelman):

You need to plug in 2 of your points to solve for "a" and for "b": \(\Large\color{black}{ 8=ab^1 }\) \(\Large\color{black}{ 512=ab^4 }\)

OpenStudy (solomonzelman):

(when you find "a" and "b", re-write the \(\Large\color{black}{ y=ab^x }\) , but with the values of "a" and "b" that you found.)